### v3

# SSM Meets Video Diffusion Models: Efficient Video Generation with Structured State Spaces

Abstract
Given the remarkable achievements in image generation through diffusion models, the research community has shown increasing interest in extending these models to video generation. Recent diffusion models for video generation have predominantly utilized attention layers to extract temporal features. However, attention layers are limited by their memory consumption, which increases quadratically with the length of the sequence. This limitation presents significant challenges when attempting to generate longer video sequences using diffusion models. To overcome this challenge, we propose leveraging state-space models (SSMs). SSMs have recently gained attention as viable alternatives due to their linear memory consumption relative to sequence length. In the experiments, we first evaluate our SSM-based model with UCF101, a standard benchmark of video generation. In addition, to investigate the potential of SSMs for longer video generation, we perform an experiment using the MineRL Navigate dataset, varying the number of frames to 64, 200, and 400. In these settings, our SSM-based model can considerably save memory consumption for longer sequences, while maintaining competitive FVD scores to the attention-based models. Our codes are available at https://github.com/shim0114/SSM-Meets-Video-Diffusion-Models.
Machine Learning, ICML
1 Introduction
Figure 1: (a) U-Net based video diffusion models consist of spatial layers and temporal layers. (b) In our temporal SSM layer, we replace an attention module with a bidirectional SSM module + a dual-layer MLP in a traditional temporal layer in VDMs. (c) Details of a bidirectional SSM in our temporal SSM layer. Following common practices, GLU is used to intermix information across different dimensions of outputs from SSM. Element-wise summation is utilized to aggregate features from bidirectional SSMs.
Expanding upon the early strides made in deep generative models, substantial research endeavors have been dedicated to the realm of video generation. Initial investigations within this domain primarily centered around frameworks including variational autoencoders (VAE) (Kingma & Welling, 2013; Kim et al., 2019; Saxena et al., 2021; Yan et al., 2021) and generative adversarial networks (GAN) (Goodfellow et al., 2014; Vondrick et al., 2016; Tulyakov et al., 2018; Ge et al., 2022). These foundational studies laid the groundwork for subsequent advancements in this field.
Research on video generation using diffusion models (Sohl-Dickstein et al., 2015; Nichol & Dhariwal, 2021; Ho et al., 2020) is cutting-edge in the field of deep generative models. The success of image generation using diffusion models, notably Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), has sparked a surge in research on applying diffusion models to video generation. This trend has been exemplified by the emergence of video diffusion models (VDMs) (Ho et al., 2022b). By harnessing the substantial representational capacity inherent in diffusion models, their application to video generation has showcased impressive performance in modeling the dynamic and intricate nature of video content (Ho et al., 2022b; Singer et al., 2022; Ho et al., 2022a)
However, research on diffusion-model-based video generation faces significant challenges in terms of computational complexity with respect to temporal sequence length. In diffusion models for video generation, attention mechanisms (Vaswani et al., 2017) are employed to capture temporal relationships (Ho et al., 2022b; Singer et al., 2022; Ho et al., 2022a; Zhou et al., 2022; Blattmann et al., 2023). Unlike recurrent models, such as LSTM layers (Chung et al., 2014; Hochreiter & Schmidhuber, 1997), attention layers allow for parallel computations, making this method widely used.In early studies on diffusion models for video generation, such as VDMs, to capture temporal relationships across video frames, temporal attention layers were added subsequent to spatial attention layers within the architecture of diffusion models for image generation, as described in Figure 1 (a). However, the memory demands of attention layers, which scale with the square of the sequence length, present substantial challenges for extending these models to handle longer sequences. Although there have been many attempts to improve the memory efficiency of attention mechanisms (Shen et al., 2018; Kitaev et al., 2020; Wang et al., 2020), these methods tend to exhibit performance deterioration compared to traditional attention mechanisms.
Recently, state-space models (SSMs) (Gu et al., 2020, 2021, 2022; Smith et al., 2023; Gu & Dao, 2023) have been identified as promising alternatives to attention mechanisms. A pioneering work by Gu et al. (2021) has enabled SSMs to capture long-term dependencies of sequential data, and their model, named the structured state space sequence model (S4), demonstrates superior performance in various benchmarks of sequence modeling. In contrast to attention mechanisms, SSMs can handle sequential data with linear complexities, so they are expected to overcome the fundamental limitation of attention-based models in many sequence modeling tasks.
Although subsequent studies have also shown that SSMs are effective across various domains (see 4.2), in the field of video generation, the application of SSMs has not been explored. This is because a methodology to effectively incorporate SSMs into video generation models, such as VDMs, has not been established. In fact, we have empirically observed that a naive approach that directly replace temporal attention layers of VDMs with SSMs works much worse than the original attention-based VDMs. To bridge this gap, in this paper, we investigate an effective approach to incorporate SSMs with video diffusion models. Our proposed model is summarized in Figure 1. A key insight of our investigation is that using bidirectional SSMs is essential to achieve good performance for SSM-based VDMs. In addition, to capture complex nonlinear dynamics of video sequences, adding a multi-layer perceptron (MLP) after the bidirectional SSM is also very effective to improve the generative performance.
In the experiments, we demonstrate that by substituting temporal attention in video diffusion models with our temporal SSM layers, it is feasible to show competitive or even better generative performance in benchmarks of video generation, e.g., UCF101 (Soomro et al., 2012), in terms of the Fréchet Video Distance (FVD) (Unterthiner et al., 2018). Moreover, we additionally observe that our SSM-based model can be successfully trained with 400-frame videos of MineRL Navigate (Guss et al., 2019; Saxena et al., 2021) with eight A100 GPUs, while attention-based VDM cannot due to huge memory consumption. We also uncover critical factors that contribute to generative qualities when utilizing SSMs for video diffusion models in ablation study.
2 Background
In this section, we provide an overview of our initial understanding of diffusion models for video generation. Diffusion models were initially introduced by Sohl-Dickstein et al. (2015), and Ho et al. (2020) further advanced the field with the introduction of denoising diffusion probabilistic models (DDPMs). DDPMs introduced a practical training algorithm, particularly well-suited for image generation. Recent advancements in diffusion models for video generation involve the extension of the DDPM architecture to accommodate video data.
2.1 Denoising Diffusion Probabilistic Models (DDPMs)
In diffusion models, the forward process involves progressively diminishing the original data signal, $\mathbf{x}_{0}$, by gradually introducing Gaussian noise as the diffusion time, $t$, advances. This sequence of transformations leads $\mathbf{x}_{0}$ to converge to pure Gaussian noise, represented as $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, at time $T$. In our study, $t$ is treated as a discrete integer within the range $[0,T]$, although some studies have considered $t$ as a continuous variable (Song et al., 2021; Kingma et al., 2021).The forward process is governed by the following following Markov process:
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$ (1)
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I}).$ (2)
In this formulation, the sequence $\sigma_{t}$ satisfies $10<\sigma_{1}<\cdots<\sigma_{T-1}<\sigma_{T}<$ 1.
The generation process in diffusion models is the reverse process. This process starts with pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$ and gradually reconstructing the data towards the original $\mathbf{x}_{0}$. During the reverse process, each step $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is modeled using a neural network.
$p_{\theta}(\mathbf{x}_{1:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),$ (3)
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\bm{\mu}_{\theta}(\mathbf{x}_{t},t),\bm{\Sigma}_{\theta}(\mathbf{x}_{t},t)),$ (4)
$p(\mathbf{x}_{T})=N(\mathbf{x}_{T};\mathbf{0},\mathbf{I}).$ (5)
Typically, $\bm{\Sigma}_{\theta}$ is set as an untrainable, time-dependent constant, $\bm{\Sigma}_{\theta}(\mathbf{x}_{t},t)=\sigma_{t}\mathbf{I}$. Additionally, with a change in the parameterization of $\bm{\mu}_{\theta}$, the reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ can be expressed as:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$
$=\mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}+\frac{\sigma_{t}^{2}}{\sqrt{\bar{\alpha}_{t}-1}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right),\sigma_{t}\mathbf{I}\right),$ (6)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\sigma_{i}^{2})$. The term $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ represents a function that predicts the noise from noisy data $\mathbf{x}_{t}$. This parameterization results in an objective function for the DDPM structured as follows:
$\mathbb{E}_{\mathbf{x}_{0},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{x_{t}},t)\right\|^{2}_{2}\right],$ (7)
where $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$.
While numerous parameterizations are recognized, such as predicting the observed data $\mathbf{x}_{0}$ from its noisy counterpart $\mathbf{x}_{t}$ or $\mathbf{v}$-prediction (Ho et al., 2022b; Salimans & Ho, 2022; Kingma et al., 2021), we chose to utilize the $\bm{\epsilon}$-prediction.
In terms of the architecture of diffusion models, 2D U-Net (Ronneberger et al., 2015) architectures are commonly used for image data. In 2D U-Net-based models, spatial attention layers are incorporated between the convolutional layers. These spatial attention layers enhance the ability to focus on relevant spatial features and improve the quality of the generated images. Although we adopted a U-Net-based architecture for the diffusion model, Peebles & Xie (2022) explored architectures based on Vision Transformers (ViT) (Dosovitskiy et al., 2020).
2.2 Architectures for Video Diffusion Models
To generate videos, diffusion models need to encapsulate both spatial and temporal features across frames. While DDPMs typically comprise a combination of U-Net and spatial attention layer, their capability is predominantly confined to spatial feature capture.
To address this limitation, Video Diffusion Models (VDMs) (Ho et al., 2022b) were introduced as an initial attempt into video generation using diffusion models. By incorporating mechanisms to capture temporal dynamics within DDPMs, VDMs enhance their capability to capture temporal features (Figure 1(a)). Temporal attention layer is commonly used in video generation diffusion models, such as VDMs, to leverage time-series dependencies. However, temporal attention requires memory proportional to the square of the sequence length, which imposes limitations on the maximum length of video sequences that can be generated at once.
In our study, we adopt VDMs as a baseline to explore the existing challenges and potential improvements in video generation using diffusion models.
3 Method
In this section, we propose the architecture of a temporal SSM (state-space model) layer for use in diffusion models for videos. Recent diffusion model-based video generation techniques capture temporal features through temporal attention layers, incurring memory costs proportional to the square of the sequence length. Recently, SSMs have emerged as a promising alternative to attention, offering linear memory costs with respect to the sequence length (Gu et al., 2020, 2021, 2022; Smith et al., 2023; Gu & Dao, 2023). We first review the recent advancements in SSMs in prior works, followed by a detailed description of our proposed temporal SSM layer architecture for video generation diffusion models.
3.1 State Space Models
Unlike the temporal attention commonly used in video diffusion models, state-space models (SSMs) enable the processing of time series with spatial complexities proportional to the sequence length. Recent studies proposed SSMs that could process inputs in parallel unlike recurrent models such as recurrent neural networks (RNNs) (Chung et al., 2014).
SSMs are widely used as sequence models that define a mapping from one-dimensional input signals $u_{k}\in\mathbb{R}$ to one-dimensional output signals $y_{k}\in\mathbb{R}$, with $s_{k}\in\mathbb{R}^{D}$ representing the hidden state. Discrete time sequences are processed as follows:
$\mathbf{s}_{k}=\mathbf{\bar{A}}\mathbf{s}_{k-1}+\mathbf{\bar{B}}u_{k},y_{k}=\mathbf{\bar{C}}\mathbf{s}_{k},$ (8)
where $\mathbf{\bar{A}}\in\mathbb{R}^{D\times D}$, $\mathbf{\bar{B}}\in\mathbb{R}^{D\times 1}$, $\mathbf{\bar{C}}\in\mathbb{R}^{1\times D}$. Unlike RNNs but akin to attention mechanisms, linear SSMs are capable of parallel computations through discrete convolutions, facilitated by the Fast Fourier Transform. The memory requirements and parallelizability of the modules managing the time series data are concisely summarized in Table 1.
Module Type Parallelizability Memory Usage
RNN × $BLD$
Attention $\surd$ $B(L^{2}+DL)$
Linear Attention $\surd$ $B(D^{2}+DL)$
SSM $\surd$ $BLD$
Table 1: Comparison of time-series handling modules in terms of parallelization capability and memory usage related to batch size ($B$), sequence length ($L$) and hidden dimension ($D$).
SSMs on their own are not a complete solution for sequence modeling. They become effective when the discrete-time matrices $\mathbf{\bar{A}},\mathbf{\bar{B}},\mathbf{\bar{C}}$ are derived from a suitable continuous-time state-space model, which enables them to efficiently manage long-range dependencies (Gu et al., 2020). In response, S4 (Gu et al., 2021) aims to learn the continuous-time SSM parameterization of $\mathbf{A},\mathbf{B},\mathbf{C}$ and the discretization rate $\Delta$, utilizing these to derive the requisite discrete-time parameters. Due to the complexity of the original transformation process, Gu et al. (2022) introduced S4D as a simplified diagonalized version of S4 that maintains similar performance in handling long sequences. We chose S4D as a backbone model for our temporal SSM layer.
As delineated in Equations 8, SSMs are inherently designed to independently handle single-input, single-output (SISO) systems. Therefore, when employing SSMs to manage multidimensional inputs, it is common practice to append a structure to capture dependencies (such as GLU (Dauphin et al., 2017)) between different dimensions of the output after the SSM (Gu et al., 2021, 2022) (see Figure 1 (c)).
3.2 Temporal SSM Layer for Diffusion Model-based Video Generation
We incorporate state-space models (SSMs) within the temporal layers for the video generation diffusion model. The structure of our proposed temporal SSM layer (Figure 1 (b)), is a key element of our study and warrants detailed discussion. This design takes cues from the structure of the temporal attention layer in VDMs, which consists of Layer Normalization (Ba et al., 2016) followed by an attention mechanism and a skip connection (He et al., 2016).
In our model, we replace the self-attention component with an SSM. We adopt a bidirectional structure, drawing from practices in Graves & Schmidhuber (2005); Wang et al. (2022); Yan et al. (2023). This choice is motivated by the inherent limitation of a single SSM, which is typically restricted to capturing unidirectional temporal transitions. By adopting a bidirectional approach, SSMs can more comprehensively understand the temporal dynamics in video data, addressing the constraints of traditional unidirectional SSMs.
Additionally, we recognize that SSM, while effective, has limitations in integrating information across different dimensions, demonstrated in Equation 8. To overcome this, we supplement the SSM with a multi-layer perceptron (MLP) after the bidirectional SSM module (Figure 1(b)). Our experiments demonstrate that this addition significantly enhances the model’s performance, proving its importance in the overall architecture.
Given an input $\mathbf{X}\in\mathbb{R}^{(B\times H\times W)\times L\times C}$, where $L$ is sequence length, $C$ is channel size, $H$ is height, $W$ is width of the input image, our proposed temporal SSM layer operates as follows:
$\mathbf{H}$ $=\text{LayerNorm}(\mathbf{X}),$
$\mathbf{F}$ $=\text{GLU}(\text{SSM}_{\text{forward}}(\mathbf{H})),$
$\mathbf{B}$ $=\text{GLU}(\text{SSM}_{\text{backward}}(\text{Flip}(\mathbf{H}))),$
$\mathbf{U}$ $=\mathbf{F}+\text{Flip}(\mathbf{B}),$
$\mathbf{O}$ $=\text{MLP}(\mathbf{U})+\mathbf{H}.$
The ‘Flip‘ function is used to reverse the temporal order of the sequence for the backward SSM and then restore the original order after processing. Element-wise summation is used to merge the forward and backward information flows, and the MLP serves to integrate across the channel dimension, which is empirically critical for performance.
4 Related Works
4.1 Deep Generative Models for Video Generation
The field of video synthesis has seen significant advancements through various studies. Prior to the emergence of diffusion models, the use of generative adversarial networks (GANs) (Goodfellow et al., 2014) dominated the scene. These methods extended traditional image-GAN frameworks to video generation, focusing on enhancing their generative capabilities (Vondrick et al., 2016; Saito et al., 2017; Tulyakov et al., 2018; Ge et al., 2022). These approaches primarily aimed to achieve their objectives by extending common architectures of GANs for image generation. Additionally, the development of long-term video generation techniques, particularly those leveraging the transitions of latent variables in variational autoencoders (VAEs) (Kingma & Welling, 2013) are also well-known. (Kim et al., 2019; Gregor & Besse, 2018; Saxena et al., 2021; Yan et al., 2021).
The advent of diffusion models in image generation (Sohl-Dickstein et al., 2015; Nichol & Dhariwal, 2021; Ho et al., 2020) marked a turning point, with their subsequent application to video distributions demonstrating promising outcomes (Ho et al., 2022b; Singer et al., 2022; Ho et al., 2022a; Harvey et al., 2022; Blattmann et al., 2023). These approaches have shown promising results. Nevertheless, the recent approaches adopt attention mechanisms to temporal layers, which requires memory proportional to the square of the sequence length, the computational and memory demands of video-diffusion models pose a substantial challenge. To mitigate this:
Spatiotemporal Downsampling  Techniques like Singer et al. (2022) and Ho et al. (2022a) have been developed to reduce computational costs by lowering spatial resolution and temporal frequency while employing temporal attention layers for capturing temporal features. These methods complement our research, which concentrates on lightening the module that captures temporal dynamics.
Latent Diffusion Models  Exploring latent diffusion models presents an alternative approach (Rombach et al., 2022), focusing on simpler latent variables rather than directly processing complex data (He et al., 2022). Some strategies involve using pre-trained image diffusion models with additional trainable layers for handling temporal features to lessen the overall computational load (Zhou et al., 2022; Blattmann et al., 2023).
Multistep Generation Schemes  In the realm of long-term video prediction using video diffusion models, multistep generation techniques are stand out (Voleti et al., 2022; Höppe et al., 2022; Harvey et al., 2022; Yin et al., 2023). These schemes generates videos through successive sampling with flexible frame conditioning, allowing for efficient long-term dependency modeling with minimal memory usage. Our research diverges from them by focusing on architectural improvements rather than sampling schemes.
4.2 Structured State Space Sequence (S4) Models
Originally introduced in Gu et al. (2021), S4 represents a sequence modeling framework that first solved all tasks in the Long-Range Arena (Tay et al., 2021). At its core lies the structured parameterization of state-space models (SSMs), offering efficient computation and demonstrating outstanding performance, in capturing long-range dependencies (Gu et al., 2020, 2021). The mathematical foundation of S4 is complex, prompting recent efforts to demystify, simplify, and improve S4 (Gu et al., 2022; Smith et al., 2023; Gu & Dao, 2023). The S4D (Gu et al., 2022) used in our study is one of such developments.
S4 and its variants have been applied across various domains, including image and video classification (Nguyen et al., 2022; Knigge et al., 2023; Islam & Bertasius, 2022; Wang et al., 2023), segmentation (Ma et al., 2024), image representation learning (Liu et al., 2024), speech generation (Goel et al., 2022), time-series generation (Zhou et al., 2023), language modeling (Mehta et al., 2023; Wang et al., 2022; Dao et al., 2023), reinforcement learning (Bar David et al., 2023; Lu et al., 2023; Deng et al., 2023). In the field of diffusion models using SSMs, DiffuSSM (Yan et al., 2023) have explored the integration of SSMs with diffusion models, replacing the computationally intensive spatial attention mechanisms in image generation with SSMs. However, the application of SSMs in video generation diffusion models remains underexplored, an area our research aims to delve into and expand upon.
Dataset UCF101 MineRL
# of Frames 16 16 64 200 400
Resolution $32\times 32$ $64\times 64$ $32\times 32$ $32\times 32$ $32\times 32$
Attention 272.152 618.005 1073.339 1032.514
Linear Attention 285.995 670.776 1175.453 1170.834 1072.001
Our SSM 226.447 634.546 1132.982 1116.339 972.306
Table 2: Comparison of FVD in UCF101 and MineRL Navigate. The absence of values for generating 400 frames using the attention mechanism is attributed to the training on NVIDIA A100 exceeding the maximum memory capacity, rendering the experiment infeasible.
5 Experiments
In this section, we introduce the incorporation of temporal SSM layers as temporal layers in diffusion models for video generation. This method facilitates the generation of longer video sequences while maintaining generative performance, overcoming the limitations imposed by memory constraints in attention-based temporal layers of video diffusion models. To empirically validate our hypothesis, we conducted a series of experiments comparing our temporal SSM layer with the temporal attention layer in video diffusion models. We also compared temporal linear attention, which has a linear spatial complexity similar to our temporal SSM layer. Additionally, we conducted an ablation study to investigate the effectiveness of each component within our temporal SSM layer. Through this analysis, we aim to identify the critical elements when integrating an SSM into the temporal layer of video diffusion models.
5.1 Experimental Setup
Datasets  We used two datasets to compare the performance of the temporal layers under varying conditions. Training details for each dataset and each frame length are shown in Appendix A. UCF101 (Soomro et al., 2012) was selected as the standard video dataset. Following Ho et al. (2022b), 13,320 videos from the training and test sets were used. We sampled clips with frame lengths of 16 and downscaled their spatial resolution to $32\times 32$ pixels or $64\times 64$ pixels for manageability. Additionally, we incorporated MineRL Navigate dataset (Guss et al., 2019; Saxena et al., 2021), which includes 961 videos for training purposes. Each video has 500 frames. This choice was motivated by our aim to validate the model’s performance with longer video sequences. We explored 64, 200, and 400 frames length, and downscaled the spatial resolution to $32\times 32$ pixels.
Baseline  We established our experimental baseline using VDMs (Ho et al., 2022b). Our analysis was meticulously designed to alter only the temporal attention layers in VDMs with our temporal SSM layers. This strategy enabled a focused examination of the impact and efficacy of the temporal SSM layers in the context of video generation, facilitating a direct comparison with the existing temporal layers.
Evaluation Metrics  In our validation process, we evaluated the sample quality of the videos generated by the trained models. To evaluate the quality of the generated videos, we employ the Fréchet Video Distance (FVD) metirc (Unterthiner et al., 2018), using an I3D network pretrained on the Kinetics-400 dataset (Carreira & Zisserman, 2017). FVD is a recognized standard for assessing the quality of generated videos (Ho et al., 2022b; Ge et al., 2022; Singer et al., 2022; Ho et al., 2022a; Harvey et al., 2022), where lower scores denote superior quality. For the UCF101 dataset, we evaluated FVD using all 13,320 videos and 10,000 generated samples. For the MineRL Navigate dataset, the calculation involves all 961 videos and 1,000 generated samples.
5.2 Benchmark Results
Figure 2: Left: Memory consumption during training with 8 NVIDIA A100 GPUs (40 GB) at a batch size of 8 and resolution of $32\times 32$. Right: Inference time for generating a sample with a single NVIDIA A100 GPU at a resolution of $32\times 32$ and $T=256$.
Figure 3: Qualitative results in MineRL Navigate (number of frames are 400). Each row represents a different sample, while each column corresponds to a sampling taken every 40 frames out of 400 frames.
We initially focus on a standard video benchmark of UCF101 to compare generative performance (measured using FVD) across different temporal layers. The quantitative results are in Table 2. The qualitative results of this generation are presented in Figure 4 (Appendix B). It is observed that the temporal SSM layer outperforms both temporal attention and temporal linear attention in generative performance in the $32\times 32$ setting and performs competitively in the $64\times 64$ setting.
SSM MLP FVD
Bidirectional # of Layers Position UCF101 MineRL-64
$\surd$ 2 post 226.447 1132.982
$\surd$ 2 pre 253.688 1129.245
2 post 669.582 1371.131
$\surd$ 1 post 267.935 1189.216
$\surd$ 0 269.811 1193.491
Table 3: Quantitative results of ablation studies in our temporal SSM layers with UCF101 16 frames and MineRL Navigate 64 frames.
We then utilize MineRL Navigate, a dataset with longer-term sequences, to observe the ability of temporal SSM layers to generate video sequences of lengths unattainable by temporal attention layers, while competitively maintaining generative performance. The transitions in these metrics as the video length increases are illustrated in Table 2. Additionally, qualitative results of the generated videos are shown in Figure 3.
In Figure 2, we present how training memory consumption and inference time vary with video sequence length for each temporal layer, using $32\times 32$ resolution images. Training memory consumption data is based on a batch size of 8, while inference times reflect sample generation on a single NVIDIA A100 GPU with the number of diffusion time steps $T$ fixed at 256. Notably, for the attention layer, experiments were capped at 275 frames instead of 400 due to memory limitations of the devices, which restrict training of attention-based video diffusion models to a maximum of approximately 275 frames.
While SSM exhibits competitive generative performance compared to temporal attention, it’s noteworthy that temporal attention encounters memory errors at sequence lengths 400, a limitation not faced by temporal SSM. This highlights SSM’s enhanced capability in handling longer sequences without compromising computational efficiency. Furthermore, SSM outperforms temporal linear attention in terms of generative performance. This finding underscores SSM’s superiority in its ability to generate high-quality videos in longer video generation.
5.3 Ablation Study of Temporal SSM Layers
In the pursuit of temporal SSM layer in video generation diffusion models, our empirical findings indicate a considerable influence of elements in the temporal layer’s architecture on model performance. We conduct an ablation study to investigate the effects of each component of temporal SSM layer (Figure 1 (b)).
Bidirectional SSM  Switching from bidirectional SSM to unidirectional SSM leads to a significant reduction in generative performance (Table 3). This result show that the bidirectional usage of SSMs significantly improves the ability of temporal SSM layer to capture temporal relationships.
MLPs in Temporal SSM Layers  In this section, we shall discuss the configuration of the MLP within our temporal SSM layer. The proposed temporal SSM layer is composed of a bidirectional SSM followed by a dual-layer MLP with GeLU activation (Hendrycks & Gimpel, 2016). The rationale behind placing the MLP after the bidirectional SSM was that the SSM part of bidirectional SSM is a single-input, single-output system that was incapable of extracting relationships with other dimensions of the input. Therefore, the MLP was used to enhance the extraction of relationships with other input dimensions.
Specifically, we compared the following configurations: placing the MLP before the bidirectional SSM rather than after, replacing the MLP with a single layer and removing the MLP altogether. The results are listed in Table 3. Transitioning from a dual-layer MLP to a single linear layer, and eventually eliminating all such components, consistently leads to a stepwise degradation in generative performance across two datasets. The results also reveal that the placement of a dual-layer MLP, whether before or after the SSM, does not significantly affect the generative capability. These findings imply that in video diffusion models employing a temporal SSM layer, the generative performance heavily relies on the expressive capacity of the temporal SSM layer, while the positioning of components relative to the SSM plays a less crucial role in determining performance.
5.4 Comparison with Prior SSM Architectures
Architecture Settings FVD
Architectures Bidirectional Dims of MLPs UCF101 MineRL-64
GSS $d_{input}\times 3$ 699.383 1683.820
BiGS $\surd$ $d_{input}\times 3$ 233.277 1157.793
BiGS $\surd$ 512 237.244 1153.669
Mamba - 669.572 1722.097
Bi-Mamba $\surd$ - 275.416 1156.891
Bi-Mamba + MLP $\surd$ 512 243.638 1138.779
Ours $\surd$ 512 226.447 1132.982
Table 4: FVD Comparison with SSM layers proposed in prior works across UCF101 16 frames and MineRL Navigate 64 frames. The resolution of both datasets is $32\times 32$ pixels. The dimensionality of the input processed by the temporal layer, $d_{\text{input}}$, is equal to the channel size, $C$, as inferred from the dimensions of the input tensor $\mathbf{X}\in\mathbb{R}^{(B\times H\times W)\times L\times C}$. In the experimental settings described in this table, the channel sizes in the upper layers of the UNet architecture follow a sequential order of 64, 128, 256, and 512, starting from the topmost layer and progressing downwards.
Recent years have seen a surge in research efforts to apply SSMs across various domains, accompanied by the proposal of numerous SSM architectures. Within this burgeoning field of study, gated state-space (GSS) architectures have emerged as one of the effective structures (Mehta et al., 2023). For instance, in Wang et al. (2022), bidirectional gated SSM (BiGS) architectures are employed for language modeling and image generation (Wang et al., 2022; Yan et al., 2023). GSS and BiGS architectures are composed of single linear layers for linear transformation, SSMs, skip connections implemented through multiplication (gating), and a two-layer MLP.
Among other effective SSM architectures, the Mamba architecture has been proposed (Gu & Dao, 2023). The Mamba architecture is characterized by a structure that alternates between 1D Convolution and SSM, and when combined with an SSM known as S6 (Gu & Dao, 2023), it has been applied to language understanding and learning of visual representations (Wang et al., 2024; Liu et al., 2024). For the learning of visual representations, a bidirectional version of Mamba, referred to here as Bi-Mamba, has also been proposed (Zhu et al., 2024). The architectures of these bidirectional SSM layers, which are more complex compared to our proposed temporal SSM layer.
It is important to note that in this section, with the goal of ensuring a fair comparison of architectures, the experiments with GSS, BiGS, Mamba, and Bi-Mamba all utilize the S4D as the SSM, similar to our temporal S4D layer. Architecture details are shown in Appendix C.
The comparisons between GSS and BiGS, as well as between Mamba and Bi-Mamba, suggest that the incorporation of bidirectionality significantly contributes to the capability of generating videos, a phenomenon not limited to our temporal SSM layers. Furthermore, introducing a Multilayer Perceptron (MLP) with a number of hidden layer dimensions that do not depend on the input size, following the SSM layer as done in our temporal SSM layers, to both BiGS and Bi-Mamba architectures, has been observed to result in performance improvements in some cases.
Then, we compare the generative performance of these structures and our proposed temporal SSM layer when used as the temporal layer in VDMs. The results indicate that, even when bidirectionality is applied and the size of the MLP’s hidden layers is aligned with ours, these models exhibit inferior generative performance compared to our simple temporal SSM layer. A potential reason for this discrepancy is that while GSS, BiGS, Mamba and Bi-Mamba focus on language and image modeling using solely the SSM layers, our study seeks a temporal SSM layer specifically tailored to capture the temporal features of video, making the former approaches excessively redundant for this purpose.
6 Discussion
Our experimental findings demonstrate that incorporating SSM into the temporal layers of diffusion models for video generation offers superior video modeling in terms of memory efficiency for handling increased sequence lengths compared to traditional models employing temporal attention, while maintaining competitive generative quality. It is empirically shown that, while attention-based video generation diffusion models struggle to generate longer videos due to memory constraints, SSM-based video diffusion models are capable of producing such videos. These outcomes underscore the adaptability of temporal SSM layers in enhancing video diffusion models and indicate their extensive potential impact on the future advancements in this domain.
This study paves the way for numerous future investigations. While we have verified the advantages of employing our temporal SSM layers within the architecture of VDMs, the application of our temporal SSM layers extends beyond this, to any model that employs temporal attention for temporal dimension. For example, adopting temporal SSM layers into the temporal components of architectures such as Make-A-Video (Singer et al., 2022) or Imagen Video (Ho et al., 2022a) could further enhance their computational efficiency by leveraging their existing spatiotemporal downsampling mechanism. Additionally, the combination of the temporal SSM layers with Latent Diffusion Models (Rombach et al., 2022) could lead to significant reductions in computational costs. Our findings also offer insights for methods that involve freezing pretrained image-generation diffusion models and training new temporal layers for video generation; simply substituting these newly inserted temporal layers with the SSM layers could be an effective strategy. Moreover, while our experiments concentrated on unconditional video generation, integrating our approach with conditional generation techniques such as classifier guidance (Dhariwal & Nichol, 2021) and classifier free guidance (Ho & Salimans, 2022) could pave the way for more efficient text-to-video (T2V) models.
The results of this study indicate that the incorporation of SSMs can lead to the development of long-term video generation models that demand fewer memory resources. This has noteworthy implications for broadening the accessibility of cutting-edge research in video generation diffusion models. Even institutions with limited computational resources can engage in this advanced field, potentially expediting the pace of research and innovation in video generation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Details in Benchmark Experiments
To ensure a fair comparison of the modules extracting the temporal relationships under the same resolution settings, all configurations except for the temporal layers, were the same in our experiments. We used NVIDIA V100 ×$\times 4$ or NVIDIA A100 ×$\times 8$ (from a cloud provider). Detailed configuration is shown in Table 5.
Dataset UCF101 MineRL
# of Frames 16 16 64 200 400
Resolution $32\times 32$ $64\times 64$ $32\times 32$ $32\times 32$ $32\times 32$
Base channel size 64 64 64 64 64
Channel multipliers 1, 2, 4, 8 1, 2, 4, 8 1, 2, 4, 8 1, 2, 4, 8 1, 2, 4, 8
Time embedding dimension 1024 1024 1024 1024 1024
Time embedding linears 2 2 2 2 2
# of attention heads (for attentions) 8 8 8 8 8
Dims of attention (for attentions) 64 64 64 64 64
SSM hidden dims (for SSMs) 512 512 512 512 512
MLP hidden dims (for SSMs) 512 512 512 512 512
Denoising timesteps (T) 256 1000 256 256 256
Loss type L2 loss of $\epsilon$ L2 loss of $\epsilon$ L2 loss of $\epsilon$ L2 loss of $\epsilon$ L2 loss of $\epsilon$
Training steps 92k 106k 174k 255k 246k
Optimizer Adam Adam Adam Adam Adam
Training learning rate 0.0003 0.0001 0.0003 0.0003 0.00005
Train batch size 64 64 8 8 8
EMA decay 0.995 0.995 0.995 0.995 0.995
GPUs V100 ×$\times 4$ A100 ×$\times 8$ V100 ×$\times 4$ A100 ×$\times 8$ A100 ×$\times 8$
Training Time 72 hours 120 hours 72 hours 100 hours 120 hours
Table 5: Experimental Setup for UCF101, MineRL Navigate datasets.
Appendix B Additional Qualitative Results
Figure 4: Qualitative generation results in UCF101 (# of frames are 16, $64\times 64$ resolution).
Appendix C Details of Architectures in Prior Works
[…]